[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jul 24, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jul 24 20:08:41 UTC 2017


HSU Shift Log for Mon Jul 24 2017 16:08:35 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: GOES16 data stopped coming in at 1015Z. Resumed at 13Z  
after talking to Mello.

AWIPS-2/NCP issues:

Other Techincal issues: Lots of missing microwave data in both the  
NRL/FNMOC pages. WindSat hasn't come in for 3 days. It is making things  
more difficult for sure. FSSE also too late to effectively use.

Miscellaneous issues: Talked with the SDM about G16 meso sectors. He says  
that often during the early daylight hours, there aren't many requests. We  
could have a sector over a tropical cyclone to help with first-daylight  
positions or invests if we coordinate ahead of time. We should probably  
consider doing this for whatever the most significant TC of the day is.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170724/954e4c7d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list